find the inverse of  f(x) = 5/2 − x, x < 2 ;

1/x, x ≥ 2


1
Expert's answer
2022-02-16T15:24:19-0500

Substitute "y" for "f(x)"


"y= \begin{cases}
 5/2-x, x<2 \\
 1/x, x\ge2
\end{cases}"

Interchange the variables "x" and "y"


"x= \begin{cases}
 5/2-y,& y<2 \\
 1/y, & y\ge2
\end{cases}"

Solve for "y"


"y= \begin{cases}
 1/x,& 0<x\le1/2 \\
 5/2-x, & x>1/2
\end{cases}"

Substitute "f^{-1}(x)" for "y"


"f^{-1}(x)= \begin{cases}
 1/x,& 0<x\le1/2 \\
 5/2-x, & x>1/2
\end{cases}"